OpenStudy (anonymous):

Write the trigonometric expression in terms of sine and cosine, and then simplify. (sin(theta)-csc(theta)) cos(theta)

OpenStudy (campbell_st):

look at the brackets 1st \[\csc(\theta) = \frac{1}{\sin(\theta)}\] \[\sin(\theta) - \csc(\theta) = \frac{\sin^2(\theta) - 1}{\sin(\theta)}\] so \[(\sin(\theta) - \csc(\theta))\cos(\theta) = (\frac{\sin^2(\theta) -1}{\sin(\theta)})\cos(\theta)\] end of of the 1st part

OpenStudy (campbell_st):

simplifying \[\sin^2(\theta) - 1 = -(1 - \sin^2(\theta)) = -\cos^2(\theta)\] so the finally \[-\cos^2(\theta) \times \frac{\cos( \theta)}{\sin(\theta)} = -\frac{\cos^3(\theta)}{\sin(\theta)}\]